One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
Answer by sofiyacherni(99)

say we call small number x
then the large number would be x+12
""The sum of the smaller number and twice the larger number is 39"" which translates into
x+2(x+12)=39
x+2x+24=39
3x+24=39
3x=15
x=5
but x was the smaller number so to find the larger number we just have to add 12
5+12=17
17 is the larger number